for a rectangular object if the lenght is 15 inches on one side and n inches on another and the height is 8 inches then what is the surface area and volume?
Answer by vihits13(20)

the volume is 15*x*8= 120x cubic inches.
the surface area is 15*x*2+ 8*x*2 + 15*8*2= 30x + 16x + 240= 240+46x square inches.